# Matching Non-Identical Objects

### v3

@@ Abstract
Not identical but similar objects are ubiquitous in our world, ranging from four-legged animals such as dogs and cats to cars of different models and flowers of various colors. This study addresses a novel task of matching such non-identical objects at the pixel level. We propose a weighting scheme of descriptors that incorporates semantic information from object detectors into existing sparse image matching methods, extending their targets from identical objects captured from different perspectives to semantically similar objects. The experiments show successful matching between non-identical objects in various cases, including in-class design variations, class discrepancy, and domain shifts (e.g., photo–drawing and image corruptions).
@@ 1 Introduction
We humans are able to find fine-grained and robust visual correspondences between two objects even when they are not necessarily identical or when they are captured in different situations. Thanks to this ability, for example, we can find out our missing cats in a gloomy backstreet by remembering them basking in the sun or assembling components to build a shelf based on their perspective drawing. In computer vision words, this needs fine-grained matching of feature points between two objects even with class discrepancy and domain shift (e.g., image corruptions).
@@
The abovementioned matching has been partially addressed in the literature. In the image matching [49, 56, 27, 17], objects to be matched are assumed identical and captured from different perspectives. This assumption also enables supervised training using dense correspondence obtained from 3D reconstruction from image pairs. Recently, Transformer-based dense matching methods have shown great performance [56, 45, 74, 23, 11, 72, 5, 67], and there are also light-weight sparse matching methods, which run in real time with a slight cost in accuracy [49, 66, 46, 38, 17, 30, 22]. In contrast, semantic correspondence models [70, 32, 7] focus on matching semantically similar parts between objects (e.g., tires of two bikes). They are trained on datasets with sparse part-to-part correspondence annotation and thus do not offer fine-grained matching.
Figure 1: Non-identical object matching by our plug-and-play module incorporated with the sparse image matching pipeline (here, SuperPoint [10] for keypoint detection and LightGlue [27] for matching). (a) design differences, (b) class discrepancy, (c) similar objects, and domain shifts such as (d) drawing v.s. photo and (e, f) different image corruptions.
@@
In this study, we generalize image matching and semantic correspondence to non-identical object matching, where objects to be matched are not necessarily identical (see Fig. 1). For example, the object pair can be street clocks of different designs or four-legged animals of different species. Further, the non-identicality naturally requires robustness, e.g., against image corruptions (e.g., blurs and brightness change) and domain shifts (e.g., photo v.s. drawing) because two objects may be captured in different environments and domains. A crucial challenge of achieving non-identical object matching is that we cannot straightforwardly resort to the supervised learning framework because of the ambiguity of non-identicality; there is no ground-truth matching between non-identical objects.
We propose a simple plug-and-play weighting module that extends off-the-shelf sparse matching methods to non-identical object matching. Given a pair of images, existing sparse matching methods first perform keypoint detection and then feed the extracted descriptors to a feature matching model to determine the matching of the keypoints. Our weighting module is inserted before the feature matching model and weights the descriptors using semantic information. Particularly, we introduce an object detector to identify objects and compute a Grad-CAM heatmap [51] to obtain weights. Previously, because of the task definition, matching methods only exploit low-level features, such as colors, edges, and texture, which are all susceptible to ambiguities and perturbations appearing in the non-identical object matching task (see Fig. 2). The low-level features are still essential to realize fine-grained matching, but the proposed module makes it robust using high-level features. Further, to avoid emphasizing irrelevant objects when an image contains more than one object, we propose Class Similarity Adapter, which determines the most relevant object pairs. We leverage the class labels from an object detector and determine object pairs based on their semantic similarity in the text embedding space.
@@
Our experiments demonstrate that our approach significantly boosts the robustness of sparse matching methods, including LightGlue [27] and GlueStick [38], in two tasks: non-identical object matching and robust image matching. In the former task, we evaluate the matching results with various image pairs (a) of the same class, (b) with class discrepancy, (c) with domain shift, and (d) of the same appearance. The sparse matchers alone result in scattered mismatching between non-corresponding areas, but combined with our method, fine-grained and consistent matching is achieved. In the latter task, we evaluate the robustness of matching on the MegaDepth-1500 [56] dataset under common corruptions [20] in terms of the AUC score of standard relative pose estimation. Our method outperforms the state-of-the-art sparse matcher under various types of image corruptions. The average AUC of our method even exceeds that of dense matchers, LoFTR [56] and Efficient LoFTR [64], which pursue matching quality at the cost of speed.
@@
  • We tackle non-identical object matching, a novel task of matching similar but unnecessarily identical objects in images, even under image corruptions and/or between different domains, which can be considered a generation of classical image matching and semantic correspodense.
  • We propose a simple plug-and-play weighting module that enhances the low-level features using semantic information and successfully extends various sparse matching methods to non-identical object matching.
  • We present a first evaluation of the robustness in image matching under common corruptions. By adding corruptions to both input images, the robustness against corruptions itself is measured. By adding them to only one of the input images, the robustness against a domain shift is measured. For both cases, our method is shown useful.
@@ 2 Related Work
Figure 2: Failure cases of non-identical object matching using the state-of-the-art matcher, LightGlue [27], and the semantic correspondence model, GeoAware-SC [70]. In semantic correspondence models, some points indicating object parts must be designated during inference to find correspondences, and automating this point selection is essential for feature point matching. Comparison of feature matching by GeoAware-SC using two methods: one that randomly extracts 50 points and the other that adaptively extracts 50 points based on the GradCAM [51] heatmap of YOLOv7 [62]. Red boxes indicate mismatches due to background or object-like textures. Orange boxes indicate semantically corresponding regions, where most points align at a part level but not at a pixel level. For example, in (d), mismatches occur between the tip of the nose of the cheetah and the nostrils of the husky.
@@
Image matching [49, 56, 27, 17] is a fundamental task in computer vision with various applications [50, 35, 8, 43, 34, 63, 57, 55] that involves finding the fine-grained correspondence between two images. Many image matching methods have two stages. A keypoint detector first extracts keypoints with their positions and local feature descriptors, and then feature matchers find correspondences between them.
Keypoint detectors.
Classically, keypoint detection was done using hand-crafted local features such as SIFT [28] and ORB [48]. Recently, learning-based methods have been developed [10, 18, 14, 61, 41, 12]. Particularly, SuperPoint [10] and DISK [59] are two popular examples. The former is based on convolutional neural networks, and the latter is based on reinforcement learning. There is also a fully differentiable and lightweight method with a minimal model configuration [14]. Unsupervised keypoint detection based on a text-to-image diffusion model [18] also exists.
Feature matchers.
Feature matchers [65, 68] find a fine-grained correspondence between images. The feature matchers are categorized into two types based on the density of matching points: sparse matchers and dense matchers. The former runs faster, while the latter gives dense and high accuracy. The focus of our study lies in sparse matchers. A seminal sparse matcher, SuperGlue [49], addresses the partial assignment problem for matching by integrating the Transformer attention mechanism [60] with optimal transport [40]. The state-of-the-art sparse matcher, LightGlue [27], improves SuperGlue by introducing a hierarchical matching structure, achieving high matching accuracy and real-time processing simultaneously. GlueStick [38] utilizes line segments for structural features. In contrast to sparse matchers, dense matchers [56, 45, 74, 23, 11, 72, 5, 67] peruse matching accuracy rather than speed. Dense matchers input an image and infer matching using all pixels as keypoints. Therefore, it does not require a keypoint detector, so it is also called a detector-free model. LoFTR [56] proposed the first transformer-based method for dense matching, which achieves highly dense matching. Recent studies [36, 71, 29] such as Efficient LoFTR [64], enabling fast and accurate matching by token aggregation, ASpanFormer [5], which controls the attention span, and ASTR [67], which focuses on the role of each pixel.
Semantic correspondence models.
Semantic Correspondence [15, 16, 33, 6, 2] is a task of identifying a semantically similar point in the target image from the given point in the source image. Specifically, it enables the correspondence of parts, such as the mouth, nose, or right eye of dogs between the source and target images. The training of a model is conducted using a specialized dataset, such as SPair-71k [33] and PF-PASCAL [15], which employ sparsely annotated part-level correspondences as supervisory data. Recently, GeoAware-SC [70] has been proposed as a model that leverages pre-trained features from two foundational models, Stable Diffusion [47] and DINOv2 [37], integrating both geometric and semantic information. This model significantly improves accuracy on major benchmarks for semantic correspondence, including the SPair-71k [33] and PF-PASCAL [15] datasets. In recent years, various models [39, 52, 69, 58, 21] have been studied. For example, CATs++ [7], which combines the local feature aggregation capacity of CNNs with the broad contextual comprehension of transformers; a model [32] employing feature maps from the reverse diffusion process in diffusion models; and unsupervised models [19] that learn semantic correspondence without requiring annotated data.
While feature matching and semantic correspondence are also the tasks of finding correspondences, our task (i.e., fine-grained matching between non-identical objects) is more challenging. First, it is hard to be addressed by the supervised learning framework because (i) manual annotation, as in semantic correspondence datasets is laborious for the pixel-level correspondence, and (ii) automatic annotation by 3D reconstruction, as in feature matching, which assumes an identical object viewed from different perspectives, cannot be done. Second, because non-identical objects generally appear with different lighting, image corruptions, or even image domains, the matching method should be designed to be robust to such diversity. As shown in Fig. 2, state-of-the-art feature matching methods and semantic correspondence models are susceptible to such variations. Several studies addressed robust matching with specific conditions, such as DarkFeat [17] for low-light and SAM [30] for viewpoint and illumination changes. We consider broader changes, including fifteen types of common corruptions [20] and several image domains (e.g., photos and sketches).
@@ 3.1 Task formulation
In the following, we present a more formal problem setup based on a standard matching pipeline. Note that it is essentially hard to completely formalize the non-identical object matching because of the ambiguity of non-identicality, so our formalization is slightly conceptual.
Problem 1 (Non-identical object matching)
Let $\mathcal{X}$} be the image domain. Given two images $\bm{x}_{A},\bm{x}_{B}\in{\mathcal{X}}$, let $D_{A}=\{(\bm{p}_{i}^{A},\bm{d}_{i}^{A})\}_{i=1}^{n_{A}}$ be $n_{A}$ keypoints in image $\bm{x}_{A}$, where $\bm{p}_{i}^{A}\in{\mathbb{R}}^{2}$ and $\bm{d}_{i}^{A}\in[0,1]^{d}$ are the position and the descriptor of $i$-th keypoint, respectively. We define $D_{B}=\{(\bm{p}_{i}^{B},\bm{d}_{i}^{B})\}_{i=1}^{n_{B}}$ for image $\bm{x}_{B}$ similarly. Let ${\mathcal{L}}(\,\cdot\,)$ be a (conceptual) matching loss function. The non-identical object matching can be formalized as follows.
@@
The matching coefficient $m_{kl}=1$ indicates that the $k$-th keypoint in image $\bm{x}_{A}$ and the $l$-th keypoint in image $\bm{x}_{B}$ is matched and otherwise $m_{kl}=0$. Each keypoint in one image has at most one matched point in the other image. This condition is implemented by the last two constraints. In the image matching task, where an identical object is captured from different angles in two images, the matching loss ${\mathcal{L}}(\{m_{kl}\}_{k,l})$ measures the discrepancy between a camera pose computed the matching $\{m_{kl}\}_{k,l}$ and the ground truth.
@@
Our task, non-identical object matching, covers a broader class of matching. Given two images $\bm{x}_{A},\bm{x}_{B}\in{\mathcal{X}}$ that contain (potentially non-identical) objects to be matched, image $\bm{x}_{I}$ can be characterized by several attributes $(o_{I},y_{I},P_{I},D_{I})$ for $I\in\{A,B\}$. If $o_{A}=o_{B}$, then the objects to be matched are identical. If $o_{A}\equiv o_{B}$, they are the same in appearance but not necessarily identical (e.g., two cars of the same model). If $y_{A}=y_{B}$, they belong to the same class (e.g., the car class). If $P_{A}=P_{B}$, then they are captured from the same camera pose. If $D_{A}=D_{B}$, they are captured in the same domain (e.g., the drawing domain).
@@
The classical image matching corresponds to the case with $o_{A}=o_{B}$ (consequently, $o_{A}\equiv o_{B}$ and $y_{A}=y_{B}$), $P_{A}\neq P_{B}$, and $D_{A}=D_{B}$. In words, it considers identical objects captured from different camera poses in the same domain. In contrast, non-identical object matching encompasses $y_{A}\neq y_{B}$ (consequently, $o_{A}\neq o_{B}$ and $o_{A}\not\equiv o_{B}$), $P_{A}\neq P_{B}$, and $D_{A}\neq D_{B}$. Note that we practically assume $o_{A}$ and $o_{B}$ to be similar in some sense. Below, we highlight two important special cases, both assuming $P_{A}\neq P_{B}$.
Figure 3: Pipeline of the matching. The keypoint detection and feature matching are done by off-the-shelf models. The proposed weighting module computes the heatmap scores of objects and weights the descriptors with this semantic information.
Class discrepancy $(y_{A}\neq y_{B})$.
@@
Four-legged animals, such as tigers and cats, are objects of different classes but have similar structures. They both have two eyes, four legs, and one tail. We humans can find a part-by-part matching between them, even with differences in color, texture, and shape. However, it is not trivial to design matching methods to find fine-grained correspondence while being robust to detail differences. If such a matching is realized, for example, one may track the running motions of unknown four-legged animals based on those of known ones.
Domain shift $(o_{A}\equiv o_{B},D_{A}\neq D_{B})$.
We humans can know the flower in the painting on the wall depicts that in the garden. This is a robust fine-grained matching of an identical object in different domains. Typical domain shifts include drawings v.s. photos and clean images v.s. corrupted images. For example, suppose that $\bm{x}_{A}$ is a photo and $\bm{x}_{B}$ is a painting. Compared to photos, drawings have fewer colors, textures, and shadows. Thus, while the descriptors of the photo $\bm{x}_{A}$ may reflect the color or texture, those of the drawing $\bm{x}_{B}$ do not, which can lead to an absent of correspondence if the feature matcher is not sufficiently robust. If robust matching between different domains is realized, for example, one can design a system that guides a worker to assemble parts from the first-person video by comparing the parts with those in a blueprint.
3.2 Technical challenges
Generalization from image matching to non-identical object matching introduces ambiguities and perturbations into matching. First, because there is no “correct matching” between two non-identical objects (e.g., a dog and a cat), one cannot resort to a simple supervised learning framework. There is no clear measure of a good matching that can be used for computing a loss to minimize. Second, domain shifts by common corruptions or image styles significantly change the low-level features. To our knowledge, existing matching methods only use low-level features such as colors, edges, and textures. As a consequence, these methods are susceptible to the domain shift, as shown in Fig. 2. The first challenge suggests that it would be better to somehow robustify pretrained image matching models instead of training a model from scratch. The second challenge indicates the necessity of introducing semantic information, which is known to be more robust against domain shifts.

### v4

# Matching Semantically Similar Non-Identical Objects

@@ Abstract
Not identical but similar objects are ubiquitous in our world, ranging from four-legged animals such as dogs and cats to cars of different models and flowers of various colors. This study addresses a novel task of matching such non-identical objects at the pixel level. We propose a weighting scheme of descriptors, Semantic Enhancement Weighting (SEW), that incorporates semantic information from object detectors into existing sparse feature matching methods, extending their targets from identical objects captured from different perspectives to semantically similar objects. The experiments show successful matching between non-identical objects in various cases, including in-class design variations, class discrepancy, and domain shifts (e.g., photo vs. drawing and image corruptions). The code is available at https://github.com/Circ-Leaf/NIOM.
@@ 1 Introduction
We humans are able to find fine-grained and robust visual correspondences between two objects even when they are not necessarily identical or when they are captured in different situations. Thanks to this ability, for example, we can find out our missing cats in a gloomy backstreet by remembering them basking in the sun or assembling components to build a shelf based on their perspective drawing. In computer vision terms, this needs fine-grained matching of feature points between two objects, even with class discrepancy and domain shift (e.g., image corruptions).
@@
The abovementioned matching has been partially addressed in the literature. In the feature matching [69, 79, 37, 24, 39], objects to be matched are assumed identical and captured from different perspectives. This assumption also enables supervised training using dense correspondence obtained from 3D reconstruction from image pairs. Recently, Transformer-based dense matching methods have shown great performance [79, 64, 103, 30, 18, 101, 8, 95, 31], and there are also light-weight sparse matching methods, which run in real time with a slight cost in accuracy [69, 92, 65, 54, 24, 43, 29, 68]. In contrast, semantic correspondence models [99, 98, 45, 34, 62] focus on matching semantically similar parts between objects (e.g., tires of two bikes). They are trained on datasets with sparse part-to-part correspondence annotation and thus do not offer fine-grained matching.
Figure 1: Non-identical object matching between various objects and image styles is achieved by our plug-and-play method with SuperPoint [16] keypoint detector and LightGlue [37] matcher.
@@
In this study, we generalize feature matching and semantic correspondence to non-identical object matching, where objects to be matched are not necessarily identical (see Fig. 1). For example, the object pair can be street clocks of different designs or four-legged animals of different species. Further, the non-identicality naturally requires robustness against image corruptions and domain shifts (e.g., photo vs. drawing) because two objects may be captured in different environments and domains. A crucial challenge of achieving non-identical object matching is that we cannot straightforwardly resort to the supervised learning framework because of the ambiguity of non-identicality; there is no ground-truth matching between non-identical objects.
We propose a plug-and-play module, Semantic Enhancement Weighting (SEW), that extends off-the-shelf matchers to non-identical object matching. The sparse matching pipeline first detects keypoints in a pair of images and feeds their descriptors into a matcher. SEW is inserted before the matcher and weights descriptors using an object detector and Grad-CAM [73]. Due to task definition, traditional matching relies on low-level features (e.g., color, edges, texture), which are prone to ambiguity and noise (see Fig. 2). While these features are still vital for fine-grained matching, our high-level feature weighting improves robustness. Further, to handle the case with multiple objects in images, we propose Non-visual Object Pairing, which determines the pairs of similar objects to be matched between images and eliminates mismatching between irrelevant objects.
@@
Our experiments demonstrate that our approach significantly boosts the robustness of sparse matching methods, including LightGlue [37] and GlueStick [54], in two tasks: non-identical object matching and robust image matching. For the former, we evaluate the matching results with various image pairs (a) of the same class, (b) with class discrepancy, (c) with domain shift, and (d) of the same appearance, and further provide a quantitative evaluation with a new annotation-free metric, Triangular Matching Consistency (TMC). While sparse matchers alone yield scattered mismatches, our method achieves fine-grained and consistent alignment, as reflected in higher TMC scores. For the latter, we evaluate the robustness of matching on the MegaDepth-1500 [79] dataset under common corruptions [27] in terms of the AUC score of standard relative pose estimation. Our method outperforms the state-of-the-art sparse matcher under various types of image corruptions. The average AUC of our method even exceeds that of dense matchers, LoFTR [79] and Efficient LoFTR [89], which pursue matching quality at the cost of speed.
@@
  • We tackle non-identical object matching, a novel task of matching similar but not necessarily identical objects in images, even under image corruptions and/or across different domains, which can be considered a generalization of feature matching and semantic correspondence.
  • We propose Semantic Enhancement Weighting (SEW), which enhances low-level features in a plug-and-play manner using semantic information and successfully extends various matchers to non-identical object matching.
  • Our experiments show that under challenging conditions such as class discrepancy, image corruptions, and domain shifts, our method yields fine-grained and consistent correspondences, validated by TMC and pose estimation.
@@ 2 Related Work
Figure 2: State-of-the-art models of sparse matching (LightGlue [37]) and semantic correspondence (GeoAware-SC [98]). The latter requires reference points in one image and finding their correspondence in the target image. To this end, 50 reference points are sampled randomly or based on the GradCAM [73] heatmap of YOLOv7 [87] (cf. Sec. D in the supplementary material). Red boxes show mismatches. Orange boxes show close matches, which are correct at the part level, but not at the pixel level. This was typical for Geo-Aware-SC (e.g., matching between the nose tip of the cheetah and the nostrils of the husky in (d)). Further, Geo-Aware-SC was about ten times slower than LightGlue. (See Tab. B in the supplementary material).
@@
Image matching [69, 79, 37, 24] is a fundamental task in computer vision with various applications [71, 51, 14, 60, 49, 88, 80, 78, 2, 93] that involves finding the fine-grained correspondence between two images. Many image matching methods have two stages. A keypoint detector first extracts keypoints with their positions and local feature descriptors, and then feature matchers find matches between them.
Keypoint detectors. Classical detectors such as SIFT [40] and ORB [67] rely on hand-crafted features. Recent learning-based methods, including SuperPoint [16] and DISK [84], achieve higher robustness, with lightweight [21] and unsupervised variants [25] also proposed.
Feature matchers. Feature matchers [90, 96] find a fine-grained correspondence between images. The feature matchers are categorized into two types based on the density of matching points: sparse matchers and dense matchers. Sparse matchers (e.g., SuperGlue [69], which combines Transformer attention with optimal transport; LightGlue [37], which introduces a hierarchical matching structure for real-time accuracy; GlueStick [54], which exploits line segments as structural features) use limited keypoints for efficiency. In contrast, dense matchers (e.g., LoFTR [79], DKM [17], ROMA [18]) pursue matching accuracy rather than speed by matching all pixels.
Semantic correspondence models. Semantic Correspondence [23, 9, 3, 45] is a task of identifying a semantically similar point in the target image from the given point in the source image, with benchmarks such as SPair-71k [48] and PF-PASCAL [22]. Recent models include GeoAware-SC [98], which integrates geometric cues with pre-trained features, and unsupervised methods without annotations [26]. It is also worth noting that semantic correspondence models run significantly slower than sparse matching methods because feature extraction depends on large generative models, such as Stable Diffusion [66]. For example, it runs roughly 10 times slower in Fig. 2.
This study addresses non-identical object matching, unifying the fine-grained alignment of feature matching and the high-level matching of semantic correspondence. This makes supervised training technically challenging but also offers unique applications (cf. Sec. 3.2). In another view, we aim for matching that runs in real-time and is robust to object identity variations, corruptions, and domain shifts. This covers broader robustness than prior studies [63, 91, 5], including DarkFeat [24] and SAM [43] for lighting changes. See Sec. H in the supplementary for detailed related work.
@@ 3.1 Task formulation
We introduce a more formal setup of non-identical object matching. To this end, we first present fine-level image matching (i.e., dense pixel-wise matching, as opposed to coarse-level image matching tasks such as semantic correspondence) as a general task and then discuss feature matching and non-identical object matching as special cases.
Problem 1 (fine-level image matching)
Let $\mathcal{X}$} be the image domain. Given two images $\bm{x}_{A},\bm{x}_{B}\in{\mathcal{X}}$, let $D_{A}=\{(\bm{p}_{i}^{A},\bm{d}_{i}^{A})\}_{i=1}^{n_{A}}$ be $n_{A}$ keypoints in image $\bm{x}_{A}$, where $\bm{p}_{i}^{A}\in[0,1]^{2}$ and $\bm{d}_{i}^{A}\in[0,1]^{d}$ are its position and descriptor of the $i$-th keypoint. We define $D_{B}=\{(\bm{p}_{i}^{B},\bm{d}_{i}^{B})\}_{i=1}^{n_{B}}$ for image $\bm{x}_{B}$ similarly. Let ${\mathcal{L}}(\,\cdot\,)$ be a (conceptual) matching loss function. The task can be formalized as follows.
@@
The matching coefficient $m_{kl}=1$ indicates that the $k$-th keypoint in image $\bm{x}_{A}$ and the $l$-th keypoint in image $\bm{x}_{B}$ is matched and otherwise $m_{kl}=0$. Each keypoint in one image has at most one matched point in the other image. This condition is implemented by the last two constraints.
In the feature matching task, where an identical object is captured from different angles in two images, matching loss ${\mathcal{L}}(\{m_{kl}\}_{k,l})$ measures the discrepancy between a camera pose computed the matching $\{m_{kl}\}_{k,l}$ and the ground truth.
@@
Our task, non-identical object matching covers a broader class of matching. Given two images $\bm{x}_{A},\bm{x}_{B}\in{\mathcal{X}}$ that contain (potentially non-identical) objects to be matched, image $\bm{x}_{I}$ can be characterized by several attributes $(o_{I},y_{I},P_{I},D_{I})$ for $I\in\{A,B\}$. If $o_{A}=o_{B}$, then the objects to be matched are identical. If $o_{A}\equiv o_{B}$, they are the same in appearance but not necessarily identical (e.g., two cars of the same model). If $y_{A}=y_{B}$, they belong to the same class (e.g., the car class). If $P_{A}=P_{B}$, then they are captured from the same camera pose. If $D_{A}=D_{B}$, they are captured in the same domain (e.g., the drawing domain).
@@
The classical feature matching corresponds to the case with $o_{A}=o_{B}$ (consequently, $o_{A}\equiv o_{B}$ and $y_{A}=y_{B}$), $P_{A}\neq P_{B}$, and $D_{A}=D_{B}$. In words, it considers identical objects captured from different camera poses in the same domain. In contrast, non-identical object matching encompasses $y_{A}\neq y_{B}$ (consequently, $o_{A}\neq o_{B}$ and $o_{A}\not\equiv o_{B}$), $P_{A}\neq P_{B}$, and $D_{A}\neq D_{B}$. Note that we practically assume $o_{A}$ and $o_{B}$ to be similar in some sense. Below, we highlight two important special cases, both assuming $P_{A}\neq P_{B}$.
Figure 3: Non-identical object warping. Homography is estimated from fine-grained matching of city clocks with different styles.
@@
Class discrepancy $(y_{A}\neq y_{B})$. Four-legged animals, such as tigers and cats, belong to different classes but have similar structures (two eyes, four legs, and a tail). Despite differences in color, texture, and shape, we humans can still find part-by-part matches. However, it is not trivial to design matching methods to find fine-grained correspondence while being robust to detail differences.
Domain shift $(o_{A}\equiv o_{B},D_{A}\neq D_{B})$. We humans can know the flower in the painting on the wall depicts that in the garden. This is a robust fine-grained matching of an identical object in different domains. Typical domain shifts include photos vs. drawings and clean vs. corrupted images. For instance, if $\bm{x}_{A}$ is a photo and $\bm{x}_{B}$ is a painting, the latter lacks colors, textures, and shadows, causing its descriptors to differ significantly. Without sufficient robustness, this discrepancy leads to missing correspondences.
3.2 Distinction from prior tasks and applications
Non-identical object matching relates to feature matching (i.e., fine-level perspective alignment) and semantic correspondence (i.e., coarse-level part correspondences between similar objects). However, achieving fine-grained and semantic alignment of different objects simultaneously is technically non-trivial and has interesting applications.
Non-identical object matching is challenging because it does not readily fit into supervised training. There is no available dataset with fine-grained human annotation of matching between distinct objects, and the construction would be laborious. In contrast, feature matching relies on supervised training with ground-truth homography from 3D reconstruction, while semantic correspondence relies on sparse part-level annotations. Yet its standard metric, PCK [48, 22], evaluates only coarse part matches and is therefore inadequate for fine-grained alignment.
The applications of non-identical object matching are unique and cannot be attained by feature matching and semantic correspondence. A fine-grained matching with robustness to class discrepancy allows transferring fine-grained landmark annotations. For example, computer vision for animals works on datasets targeting a single animal species, such as facial datasets of dogs [46], cats [47], and cattle [12]. Non-identical object matching allows transferring their annotations to other animals. Its robustness to domain shift can also compare the sketch or blueprint of a product to a real one, which helps workers assemble parts and build it up. Figure 3 illustrates a simple example, non-identical object warping, that is not covered by prior tasks.
Figure 4: Pipeline of the matching. The keypoint detection and feature matching are done by off-the-shelf models. The proposed plug-and-play module, SEW, computes the heatmap scores of objects and weights the descriptors with this semantic information.